A bank's assets are $90 million, and its liabilities are $71 million. Its assets increase by 10 percent, and its liabilities increase by 6 percent. What is
the percentage change in the bank's capital, or net worth?
1
Expert's answer
2020-04-07T08:19:14-0400

Asset=$90 m

liabilities=$71 m

"\text{new assets}= (110*90)/100" =$99m

"\text{new liabilities}=(106*71)/100" =$75.26

"capital=assets\u2212liabilities"

"\text{initial capital}=90-71" =$19m

"\text{new capital}=99-75.26" =$23.74m

"\text{change in capital}=23.74-19" =$4.74m

Percentage change in capital,%c therefore becomes;

%c=(4.74/19)*100=24.95%